# AI-enhanced Collective Intelligence: The State of the Art and Prospects

### v2

@@ 2.6.1. Perception and reaction
Humans tend to be more accepting of AI when it makes mistakes. Some studies find that users have a more favorable impression of the imperfect robot than the perfect robot when the robot behaves adequately after making mistakes (Yasuda and Matsumoto, 2013). When both human drivers and AI make errors in cases of human–machine shared-control vehicles, the blame attributed to the machine is reduced (Awad et al., 2020). Research has found that errors occasionally performed by a humanoid robot can increase its perceived human likeness and likability (Salem et al., 2013). A recent study shows humans accept their AI teammate’s decision less often when they are deceived about the identity of the AI as another human (Zhang et al., 2023a). The nature of the task requested by the robot, e.g., whether its effects are revocable as opposed to irrevocable, significantly impacts participants’ willingness to follow its instructions (Salem et al., 2015). When taking AI’s advice, humans must know the AI system’s error boundary and decide when to accept or override AI’s recommendation (Bansal et al., 2019).
2.6.2. Trust
@@
In AI-assisted decision-making, where the individual strengths of humans and AI optimize the joint decision outcome, the key to success is appropriately calibrating human trust (Zhang et al., 2020). Trust between humans and AI plays a pivotal role in realizing collective intelligence (Gupta et al., 2023). Previous research highlights the necessity of understanding factors that enable or hinder the formation, maintenance, and repair of trust in human-AI collaborations (Gupta et al., 2023). Key factors influencing this trust include the perceived competence, benevolence, and integrity (McAllister, 1995) of the AI systems, mirroring the trust dynamics in human relationships.
@@
The level of anthropomorphism of machines can be a predictor of the trust level of the human participants (Natarajan and Gombolay, 2020). Studies indicate that while anthropomorphic features in AI agents initially create positive impressions, this effect is often short-lived (Gupta et al., 2023). Trust tends to decline, especially when the agent’s capabilities are not clearly presented, and its performance fails to meet users’ expectations (Glikson and Woolley, 2020).
@@
Trust in a teammate may evolve over time irrespective of team performance (McNeese et al., 2021a). Knowing when to trust the AI allows the human expert to appropriately apply their knowledge, improving decision outcomes in cases where the model is likely to perform poorly (Zhang et al., 2020). A previous study also proposes a framework highlighting trust relationships within human-AI teams. It acknowledges the multilevel nature of team trust, which considers individual, dyads, and team levels as a whole (Ulfert et al., 2023). Another study points out that the trust level of a team needs to be uniform, as uncooperative members could undermine the team’s ability to reach intended goals (De Jong et al., 2021).
2.6.3. Explainability
@@
In human-AI decision-making scenarios, people’s awareness of how AI works and its outcomes is critical in building a relationship with the system. AI explanation can support people in justifying their decisions (Ferreira and Monteiro, 2021). A study on content moderation found that when explanations were provided for why content was taken down, removal decisions made jointly by humans and AI were perceived as more trustworthy and acceptable than the same decisions made by humans alone (Wang and Kim, 2023). Many researchers have proposed using explainable AI (XAI) to enable humans to rely on AI advice appropriately and reach complementary team performance (Hemmer et al., 2022).
@@
Previous research finds that AI explanations can increase human acceptance of AI recommendations, regardless of their correctness (Bansal et al., 2021), leading to an over-reliance, which threatens the performance of human-AI decision-making. Whereas another study finds that there are scenarios where AI explanations can reduce over-reliance (Vasconcelos et al., 2023) using a cost-benefit framework. Some studies show that cognitive forcing can reduce over-reliance compared to the simple explainable AI approaches (Buçinca et al., 2021), with a trade-off that the designs most effective in reducing over-reliance were the least favored by users (Buçinca et al., 2021). The impact of AI explanations on decision-making tasks is contingent upon individuals’ diverse levels of domain expertise (Wang and Yin, 2021).
3. Applications and implications
@@
The AI-enhanced collective intelligence approach is increasingly applied across various domains, offering innovative solutions to challenges such as community response to climate change, environment, and sustainability challenges (Nesta, 2024), etc. AI serves as a vital decision support tool for policymakers in detecting misinformation (Kou et al., 2022) and real-time crisis management, exemplified during the COVID-19 pandemic. Its application extends to high-stake areas, including medical diagnosis (Lebovitz et al., 2022) and criminal justice (Rigano, 2019). AI’s integration into collective intelligence efforts enables significant amplification of its impact and scalability (Verhulst, 2018).
@@
Figure 2. Distribution of AI-enhanced collective intelligence cases by application area based on dataset curated by Supermind Design (Design, 2024).
@@
Here, we present our analysis utilizing the Supermind Design database (Design, [n. d.]), which includes 857 categorized cases of AI-enhanced collective intelligence by application areas. Notably, while some cases might not directly involve AI or collective intelligence depending on the exact definitions, the database collectively offers a comprehensive overview. It provides insights into the state of the art in AI-CI applications. As shown in Fig. 2, the majority of applications, approximately twenty percent, are found in the public sector and NGOs, followed by high-tech, media, telecommunications, entertainment, and hospitality. Fewer instances are found in supply chain, real estate, and agriculture. It is important to note that several cases classified under ”Public sector, NGO” are private sector initiatives addressing public issues. We will analyze and highlight specific AI-CI examples across these diverse domains in the upcoming sections.
3.1. Public sector, NGO
@@
Safecity (Safecity, [n. d.]) is a platform that crowdsources personal stories of sexual harassment and abuse in public spaces. It collaborates with the AI partner OMDENA (Omdena, [n. d.]) to build solutions to help women at risk of sexual abuse by predicting places at a high risk of sexual harassment incidents utilizing convolutional neural networks and LSTM (Long-Short-Term-Memory).
@@
Ushahidi (Ushahidi, [n. d.]) is a non-profit online platform for crowdsourcing information, mapping, and visualizing data to respond to and inform decisions on social issues and crises. It aggregates crowdsourced information and reports from individuals across the globe. Ushahidi is developing new tools to enhance data management and analysis, such as leveraging machine learning techniques to improve the efficiency of processing crisis reports, thus enabling faster response time (Nesta, [n. d.]).
@@
While NGOs and the public sector might seem like fertile grounds for AI-CI application, it is crucial to consider the potential high risk issues that may arise and their impact during the design and deployment of such AI-CI systems. A notable instance of a data-driven tool for social good that diverged from its original accuracy and effect is the infamous case of Google Flu (Lazer et al., 2014).
3.2. High Tech (software)
@@
Bluenove (Bluenove, 2023), a technology and consulting company that pioneers massive collective intelligence, focuses on mobilizing communities on a large scale. Their Assembl (Bluenove, 2024) platform facilitates open conversations among participants, employing advanced AI technologies such as natural language recognition, semantic analysis, and emotion analysis to categorize content published on the platform automatically.
@@
Figure Eight (Eight, [n. d.]) (formerly known as CrowdFlower (CrowdFlower, 2024)) was a human-in-the-loop machine learning and AI company acquired by Appen (Appen, 2024) in 2019. It leverages crowdsourced human intelligence to perform tasks such as text transcription or image annotation to train machine learning algorithms, which can be used in various applications.
@@
The recent advancements in high-tech technologies, especially the burgeoning field of LLM software, hold the potential for revolutionary developments. Nevertheless, the high risk associated with errors in automated or AI-generated software (Zhang et al., 2023b) necessitates the inclusion of human intelligence in these projects.
3.3. Media, telecommunication, entertainment, hospitality
@@
Civil War Photo Sleuth (Sleuth, [n. d.]) combines facial recognition technology and community effort to uncover lost identities in photographs from the American Civil War era. It leverages a dedicated community with a keen interest in the US Civil War and focuses on more accurately tagging and identifying individuals in historical photographs.
@@
Bellingcat (Bellingcat, 2024) is an independent investigative online journalism community specializing in fact-checking and open-source intelligence. It brings together a collective of researchers, investigators, and citizen journalists to publish investigation findings. Utilizing various AI tools in investigative work, Bellingcat obtains insights from massive open-source data, such as satellite imagery, photographs, video recordings, and social media posts.
@@
Facing the emerging challenge of deepfake videos, research indicates that a hybrid system combining human judgment with AI models outperforms either humans or models alone in detecting deepfakes (Groh et al., 2022), underscoring the value of integrating human perceptual skills with AI technology in addressing the complexities posed by deepfake content.
3.4. Energy, natural resources
@@
Litterati (de Leon, 2019) is a mobile app and community platform that combats litter by empowering people to take action. The app allows the worldwide community to contribute their observations and findings through photographing and documenting occurrences, creating a crowdsourced database of litter data. It also provides a platform for individuals to connect, share, and support each other’s initiatives. The platform uses AI, such as machine learning and computer vision techniques, to automatically recognize different objects and materials and efficiently organize the collected litter data. Similarly, OpenLitterMap (OpenLitterMap, 2023) is another example of making a clearer planet using AI-enhanced collective intelligence.
@@
eBird (Sullivan et al., 2014) is a mobile app and community platform for biodiversity and supports conservation initiatives to protect bird species and their habitats. It harnesses the power of a global community of birdwatchers and citizen scientists to collectively report bird observations and contribute to a shared knowledge base. eBird uses computer vision techniques to identify bird species from submitted photographs and recordings and employs machine learning predictive models to forecast bird migration patterns and species distributions.
@@
Tackling environmental challenges is crucial, and large-scale, crowd-based AI-enhanced collective intelligence projects play a significant role in this endeavor. Nevertheless, integrating AI with human efforts might dehumanize the work and demotivate human participants (Trouille et al., 2019). This balance between efficiently leveraging AI and maintaining human engagement is a key consideration in such projects.
3.5. Education and academia
@@
Kialo Edu (Kialo, 2024) is an argument mapping and debate site dedicated to facilitating collaborative and critical thinking. Kialo uses natural language processing techniques to analyze arguments, cluster topics, and identify discussion patterns. It allows users to participate in structured debates and discussions, promoting constructive engagement among students. This interaction helps them deeply understand and analyze the core aspects of the topics under discussion.
@@
Zooniverse (Zooniverse, 2024) is a platform for people-powered research where volunteers assist professional researchers. Galaxy Zoo is a prominent project within the Zooniverse platform. This crowdsourced astronomy project invites people to assist in the morphological classification of large numbers of galaxies. It uses AI algorithms to learn patterns and characteristics indicative of different galaxy types by training on large datasets previously classified by citizen scientists. The combination of efforts from both the crowd and AI enables Galaxy Zoo to harness the power of human expertise and the efficiency of AI to scale up the analysis and understanding of galaxies.
@@
A study analyzing the contributor demographics of the Zooniverse project revealed uneven geographical distribution and a gender imbalance, with approximately 30% of the citizen scientists being female (Ibrahim et al., 2021). This finding highlights concerns about representativeness and diversity in citizen science initiatives, indicating potential biases in the data and insights from such projects. AI can be used to train human contributors (Zevin et al., 2017), to de-bias collective intelligence, and to evaluate and combine human solutions (Kleinberg et al., 2018).
3.6. Healthcare
@@
The Human Diagnosis Project (Human Dx) (Dx, 2024) is an open online system of collective medical insights that provides healthcare support to patients. Human Dx enables healthcare professionals worldwide to contribute expertise and collaborate on complex diagnostic cases (Bresnick, 2017). The platform facilitates idea exchange and mutual learning to ultimately benefit patient care by utilizing machine learning techniques that automatically learn from classifying patterns in crowdsourced data. Human Dx makes the most of limited medical resources by harnessing the collective intelligence of human physicians and AI, enabling more accurate, affordable, and accessible care for those in need.
@@
CrowdEEG (CrowdEEG, 2024) is a collaborative annotation tool for medical time series data. The project combines human and machine intelligence for scalable and accurate human clinical EEG data analysis. It trains machine learning algorithms using feedback from clinical experts and non-expert crowds to perform feature detection or classification tasks on medical time series data.
@@
In healthcare, the use of AI-CI poses the risk of amplifying existing biases in human-only diagnosis once we combine AI with human judgment (Groh et al., 2024). Consequently, errors resulting from such biased decision-making can erode patients’ trust not only in the AI systems but also in the medical professionals utilizing them.
3.7. Financial services
@@
Numerai (NUMERAI, 2024) is an AI-run quant hedge fund built on crowdsourced machine learning models. Numerai hosts an innovative data science tournament in which a global network of data scientists develops machine-learning models to predict stock markets. Based on the combined knowledge of the participants, Numerai combines and aggregates these crowdsourced predictive models into an ensemble model to derive investment strategies in the financial markets.
@@
CryptoSwarm AI (AI, 2024) is a forecasting service that provides rigorous insights and intelligence on cryptocurrencies and other Web3 assets. A combination of AI technology and real-time human insights powers it. CryptoSwarm AI uses Swarm AI (AI, [n. d.]) to amplify the collective intelligence of the online communities quickly, enabling networked groups to converge on optimized solutions in real time.
@@
In financial markets, research generally indicates positive impacts on market efficiency due to the superior information processing and optimization capabilities of machines, coupled with human anticipation of these effects. However, this improvement is sometimes disrupted by rapid price spikes and crashes, resulting from the herding behavior of machines and the challenges humans face in intervening at extremely fast timescales (Tsvetkova et al., 2024).
3.8. Supply chain, real estate
@@
MarineTraffic (MarineTraffic, 2024a) is a web-based platform that creates a global network of vessel tracking information, enabling users to track vessels, monitor maritime traffic, analyze trends, and make informed decisions. It relies on the collective contributions of the maritime community, who voluntarily transmit real-time information about vessel positions and movements from various sources, including AIS (automatic identification system) data and satellite data. MarineTraffic employs neural network architectures for automatic maritime object detection using satellite imagery (Bereta et al., 2020). More examples of AI applications are available at MarineTraffic Research (MarineTraffic, 2024b).
@@
Waze (Waze, [n. d.]) is a GPS navigation software that works on smartphones and other computers. It analyzes crowdsourced GPS data and user-submitted information along the route in real-time and uses community editing to ensure the accuracy of the map data. Waze uses AI algorithms to predict traffic, optimize routes (Link, 2019), and provide personalized recommendations by analyzing driver behaviors. Waze has recently implemented integration with Google Assistant, enabling users to use the application by utilizing the voice command ”Hey Google” (Google, [n. d.]).
@@
By improving efficiency, reducing waste, and enabling more equitable distribution and accessibility of resources and spaces, AI can play a significant role in addressing inequalities in access to resources, paving the way for sustainable and efficient infrastructure development (Cowls et al., 2021).
3.9. Agriculture
@@
WeFarm (Nesta, 2019) is a social networking platform connecting the small-scale farming community. It helps millions of African farmers meet, exchange solutions to their questions, and trade equipment and supplies. Farmers ask each other questions about agriculture and promptly receive content and ideas from other farmers worldwide through crowdsourcing. Natural language processing processes and understands the text-based messages farmers send in different regions. Machine-learning matching algorithms consider farmers’ needs and expertise, identify the most suitable responses, and present them to the farmer who posed the question.
@@
Mercy Corps AgriFin (AgriFin, 2023) harnesses the power of AI and a global network of partners to transform agriculture for smallholder farmers. By utilizing state-of-the-art imagery, modeling, and analysis, AI enables precise management of crops and adaptation strategies against climate change impacts. AgriFin’s integration of digital technology and data, supported by a global network, empowers these farmers in a digitally interconnected world.
@@
In light of the global food crisis and the quest for sustainable solutions, the importance of AI-enhanced collective intelligence platforms becomes increasingly critical. These platforms can significantly contribute to addressing agricultural challenges and transitioning towards global sustainable food systems (Marvin et al., 2022).
4. Challenges and outlooks
@@
The application areas of AI-enhanced collective intelligence continually expand as technology advances and new opportunities emerge. As introduced in the previous section, successful examples of AI-enhanced collective intelligence exist in various domains. However, challenges coexist with opportunities. When realizing human-AI collective intelligence, many aspects must be considered, such as human-machine communication, trust, crowd retention, technology design, and ethical issues (Tsvetkova et al., 2017b). Here, we briefly discuss some of these challenges and call for further work on each aspect.
4.1. Communication
@@
Clear and effective communication between humans and AI is the engine for collective intelligence to emerge (Gupta et al., 2023). Communication refers to the process of exchanging information between teammates (Salas et al., 2009). It is essential for team performance as it contributes to the development and maintenance of shared mental models and the successful execution of necessary team processes (Salas et al., 2009; Stowers et al., 2021).

### v3

@@ Introduction
Humans tend to be more accepting of AI when it makes mistakes. Some studies find that users have a more favorable impression of the imperfect robot than the perfect robot when the robot behaves adequately after making mistakes 161. When both human drivers and AI make errors in cases of human-machine shared-control vehicles, the blame attributed to the machine is reduced 162. Research has found that errors occasionally performed by a humanoid robot can increase its perceived human likeness and likability 163. A recent study shows humans accept their AI teammate’s decision less often when they are deceived about the identity of the AI as another human 164. The nature of the task requested by the robot, e.g., whether its effects are revocable as opposed to irrevocable, significantly impacts participants’ willingness to follow its instructions 165. When taking AI’s advice, humans must know the AI system’s error boundary and decide when to accept or override AI’s recommendation 166.
Trust
@@
In AI-assisted decision-making, where the individual strengths of humans and AI optimize the joint decision outcome, the key to success is appropriately calibrating human trust 167. Trust between humans and AI plays a pivotal role in realizing collective intelligence 168. Previous research highlights the necessity of understanding factors that enable or hinder the formation, maintenance, and repair of trust in human-AI collaborations 168. Key factors influencing this trust include the perceived competence, benevolence, and integrity 169 of the AI systems, mirroring the trust dynamics in human relationships.
@@
The level of anthropomorphism of machines can be a predictor of the trust level of the human participants 170. Studies indicate that while anthropomorphic features in AI agents initially create positive impressions, this effect is often short-lived 168. Trust tends to decline, especially when the agent’s capabilities are not clearly presented and its performance fails to meet users’ expectations 171.
@@
Trust in a teammate may evolve over time irrespective of team performance 172. Knowing when to trust the AI allows the human expert to appropriately apply their knowledge, improving decision outcomes in cases where the model is likely to perform poorly 167. A previous study also proposes a framework highlighting trust relationships within human-AI teams. It acknowledges the multilevel nature of team trust, which considers individual, dyads, and team levels as a whole 173. Another study points out that the trust level of a team needs to be uniform, as uncooperative members could undermine the team’s ability to reach intended goals 174.
Explainability
@@
In human-AI decision-making scenarios, people’s awareness of how AI works and its outcomes is critical in building a relationship with the system. AI explanation can support people in justifying their decisions 175. A study on content moderation found that when explanations were provided for why the content was taken down, removal decisions made jointly by humans and AI were perceived as more trustworthy and acceptable than the same decisions made by humans alone 176. Many researchers have proposed using explainable AI (XAI) to enable humans to rely on AI advice appropriately and reach complementary team performance 39. However, some studies suggest that XAI can be associated with a white-box paradox 177, potentially leading to null or detrimental effects.
@@
Previous research finds that AI explanations can increase human acceptance of AI recommendations, regardless of their correctness 178, leading to an over-reliance, threatening the performance of human-AI decision-making. Whereas another study finds that there are scenarios where AI explanations can reduce over-reliance 179 using a cost-benefit framework. The impact of AI explanations on decision-making tasks is contingent upon individuals’ diverse levels of domain expertise 180. Additionally, some studies advocate for the concept of causability 181, which measures whether and to what extent humans can understand a given machine explanation to develop effective human-AI interfaces, particularly in medical AI.
Applications and implications
@@
The AI-enhanced collective intelligence approach is increasingly applied across various domains, offering innovative solutions to challenges such as community response to climate change, environment, and sustainability challenges 182, etc. AI serves as a vital decision support tool for policymakers in detecting misinformation 183 and real-time crisis management, exemplified during the COVID-19 pandemic. Its application extends to high-stake areas, including medical diagnosis 184 and criminal justice 185. AI’s integration into collective intelligence efforts enables significant amplification of its impact and scalability 186.
@@
Figure 2: Distribution of AI-enhanced collective intelligence cases by application area based on dataset curated by Supermind Design 187.
@@
Here, we present our analysis utilizing the Supermind Design database 188, which includes over one thousand actual examples of AI-enhanced collective intelligence, with 938 categorized into twelve application areas. Notably, while some cases might not directly involve AI or collective intelligence depending on the exact definitions, the database collectively offers a comprehensive overview. It provides insights into the state-of-the-art in AI-CI applications. As shown in Fig. 2, the majority of applications, approximately twenty percent, are found in the public sector and NGOs, followed by high-tech, media, telecommunications, entertainment, and hospitality. Fewer instances are found in supply chain, real estate, and agriculture. It is important to note that several cases classified under “Public sector, NGO” are private sector initiatives addressing public issues. In the upcoming sections, we will analyze and highlight specific AI-CI examples across these diverse domains. The supplemental information Table S2 provides a summary of the CI and AI aspects of these examples.
Public sector, NGO
@@
Safecity 189 is a platform that crowdsources personal stories of sexual harassment and abuse in public spaces. It collaborates with the AI partner OMDENA 190 to build solutions to help women at risk of sexual abuse by predicting places at a high risk of sexual harassment incidents utilizing convolutional neural networks and LSTM (Long-Short-Term-Memory).
@@
Ushahidi 191 is a non-profit online platform for crowdsourcing information, mapping, and visualizing data to respond to and inform decisions on social issues and crises. It aggregates crowdsourced information and reports from individuals across the globe. Ushahidi is developing new tools to enhance data management and analysis, such as leveraging machine learning techniques to improve the efficiency of processing crisis reports, thus enabling faster response time 192.
@@
While NGOs and the public sector might seem like fertile grounds for AI-CI applications, it is crucial to consider the potential high-risk issues that may arise and their impact during the design and deployment of such AI-CI systems. A notable instance of a data-driven tool for social good that diverged from its original accuracy and effect is the infamous case of Google Flu 193.
Just as AI-enhanced collective intelligence helps address social issues and crises in the public sector and NGOs, it also drives innovation and efficiency in the high-tech industry.
High Tech (software)
@@
Bluenove 194, a technology and consulting company that pioneers massive collective intelligence, focuses on mobilizing communities on a large scale. Their Assembl 195 platform facilitates open conversations among participants, employing advanced AI technologies such as natural language recognition, semantic analysis, and emotion analysis to categorize content published on the platform automatically.
@@
Figure Eight 196 (formerly known as CrowdFlower 197) was a human-in-the-loop machine learning and AI company acquired by Appen 198 in 2019. It leverages crowdsourced human intelligence to perform tasks such as text transcription or image annotation to train machine learning algorithms, which can be used in various applications.
@@
The recent advancements in high-tech technologies, especially the burgeoning field of LLM software, hold the potential for revolutionary developments. Nevertheless, the high risk associated with errors in automated or AI-generated software 199 necessitates the inclusion of human intelligence in these projects.
Media, telecommunication, entertainment, hospitality
Similar to high tech, more traditional information and communication technologies, such as the media sector, harness AI-enhanced CI for purposes such as investigative journalism, historical research, and combating misinformation.
@@
Civil War Photo Sleuth 200 combines facial recognition technology and community effort to uncover lost identities in photographs from the American Civil War era. It leverages a dedicated community with a keen interest in the US Civil War and focuses on more accurately tagging and identifying individuals in historical photographs.
@@
Bellingcat 201 is an independent investigative online journalism community specializing in fact-checking and open-source intelligence. It brings together a collective of researchers, investigators, and citizen journalists to publish investigation findings. Utilizing various AI tools in investigative work, Bellingcat obtains insights from massive open-source data, such as satellite imagery, photographs, video recordings, and social media posts.
@@
Facing the emerging challenge of deepfake videos, research indicates that a hybrid system combining human judgment with AI models outperforms either humans or models alone in detecting deepfakes 202, underscoring the value of integrating human perceptual skills with AI technology in addressing the complexities posed by deepfake content.
Energy, natural resources
@@
Swithing to a more industrial environment, in the energy and natural resources sector, collective intelligence combined with AI can play a role in addressing environmental challenges through community-driven data collection and analysis. Litterati 203 is a mobile app and community platform that combats litter by empowering people to take action. The app allows the worldwide community to contribute their observations and findings through photographing and documenting occurrences, creating a crowdsourced database of litter data. It also provides a platform for individuals to connect, share, and support each other’s initiatives. The platform uses AI, such as machine learning and computer vision techniques, to automatically recognize different objects and materials and efficiently organize the collected litter data. Similarly, OpenLitterMap 204 is another example of making a clearer planet using AI-enhanced collective intelligence.
@@
eBird 205 is a mobile app and community platform for biodiversity and supports conservation initiatives to protect bird species and their habitats. It harnesses the power of a global community of birdwatchers and citizen scientists to collectively report bird observations and contribute to a shared knowledge base. eBird uses computer vision techniques to identify bird species from submitted photographs and recordings and employs machine learning predictive models to forecast bird migration patterns and species distributions.
@@
Tackling environmental challenges is crucial, and large-scale, crowd-based AI-enhanced collective intelligence projects play a significant role in this endeavor. Nevertheless, integrating AI with human efforts might dehumanize the work and demotivate human participants 206. This balance between efficiently leveraging AI and maintaining human engagement is a key consideration in such projects.
Education and academia
@@
Education and academia is another area that benefit from AI-enhanced collective intelligence by fostering collaborative learning environments and supporting large-scale research projects. Kialo Edu 207 is an argument mapping and debate site dedicated to facilitating collaborative and critical thinking. Kialo uses natural language processing techniques to analyze arguments, cluster topics, and identify discussion patterns. It allows users to participate in structured debates and discussions, promoting constructive engagement among students. This interaction helps them deeply understand and analyze the core aspects of the topics under discussion.
@@
Zooniverse 208 is a platform for people-powered research where volunteers assist professional researchers. Galaxy Zoo is a prominent project within the Zooniverse platform. This crowdsourced astronomy project invites people to assist in the morphological classification of large numbers of galaxies. It uses AI algorithms to learn patterns and characteristics indicative of different galaxy types by training on large datasets previously classified by citizen scientists. The combination of efforts from both the crowd and AI enables Galaxy Zoo to harness the power of human expertise and the efficiency of AI to scale up the analysis and understanding of galaxies.
@@
A study analyzing the contributor demographics of the Zooniverse project revealed uneven geographical distribution and a gender imbalance, with approximately 30% of the citizen scientists being female 4. This finding highlights concerns about representativeness and diversity in citizen science initiatives, indicating potential biases in the data and insights from such projects. AI can be used to train human contributors 209, to de-bias collective intelligence, and to evaluate and combine human solutions 210.
Healthcare
@@
In a similar vein, the healthcare sector also leverages AI-enhanced collective intelligence, addressing challenges such as diagnostic accuracy and medical data analysis, which are critical for improving patient outcomes and advancing medical research. The Human Diagnosis Project (Human Dx) 211 is an open online system of collective medical insights that provides healthcare support to patients. Human Dx enables healthcare professionals worldwide to contribute expertise and collaborate on complex diagnostic cases 212. The platform facilitates idea exchange and mutual learning to ultimately benefit patient care by utilizing machine learning techniques that automatically learn from classifying patterns in crowdsourced data. Human Dx makes the most of limited medical resources by harnessing the collective intelligence of human physicians and AI, enabling more accurate, affordable, and accessible care for those in need.
@@
CrowdEEG 213 is a collaborative annotation tool for medical time series data. The project combines human and machine intelligence for scalable and accurate human clinical EEG data analysis. It trains machine learning algorithms using feedback from clinical experts and non-expert crowds to perform feature detection or classification tasks on medical time series data.
@@
In healthcare, the use of AI-CI poses the risk of amplifying existing biases in human-only diagnosis once we combine AI with human judgment 214. Consequently, errors resulting from such biased decision-making can erode patients’ trust not only in the AI systems but also in the medical professionals utilizing them.
Financial services
@@
AI-enhanced collective intelligence in financial services optimizes investment strategies and improve market predictions, demonstrating the powerful synergy between human expertise and machine learning. Numerai 215 is an AI-run quant hedge fund built on crowdsourced machine learning models. Numerai hosts an innovative data science tournament in which a global network of data scientists develops machine-learning models to predict stock markets. Based on the combined knowledge of the participants, Numerai combines and aggregates these crowdsourced predictive models into an ensemble model to derive investment strategies in the financial markets.
@@
CryptoSwarm AI 216 is a forecasting service that provides rigorous insights and intelligence on cryptocurrencies and other Web3 assets. A combination of AI technology and real-time human insights powers it. CryptoSwarm AI uses Swarm AI 217 to amplify the collective intelligence of the online communities quickly, enabling networked groups to converge on optimized solutions in real time.
@@
In financial markets, research generally indicates positive impacts on market efficiency due to the superior information processing and optimization capabilities of machines, coupled with human anticipation of these effects. However, this improvement is sometimes disrupted by rapid price spikes and crashes, resulting from the herding behavior of machines and the challenges humans face in intervening at extremely fast timescales 218.
Supply chain, real estate
@@
The supply chain and real estate sectors also benefit from AI-enhanced collective intelligence, improving efficiency, transparency, and decision-making through real-time data integration and analysis. MarineTraffic 219 is a web-based platform that creates a global network of vessel tracking information, enabling users to track vessels, monitor maritime traffic, analyze trends, and make informed decisions. It relies on the collective contributions of the maritime community, who voluntarily transmit real-time information about vessel positions and movements from various sources, including AIS (automatic identification system) data and satellite data. MarineTraffic employs neural network architectures for automatic maritime object detection using satellite imagery 220. More examples of AI applications are available at MarineTraffic Research 221.
@@
Waze 222 is a GPS navigation software that works on smartphones and other computers. It analyzes crowdsourced GPS data and user-submitted information along the route in real-time and uses community editing to ensure the accuracy of the map data. Waze uses AI algorithms to predict traffic, optimize routes 223, and provide personalized recommendations by analyzing driver behaviors. Waze has implemented integration with Google Assistant, enabling users to use the application by utilizing the voice command “Hey Google” 224.
@@
By improving efficiency, reducing waste, and enabling more equitable distribution and accessibility of resources and spaces, AI can play a significant role in addressing inequalities in access to resources, paving the way for sustainable and efficient infrastructure development 225.
Agriculture
@@
Finally, the agriculture sector leverages AI-enhanced collective intelligence to improve farming practices, resource management, and crop yields, addressing global food security challenges. WeFarm 226 is a social networking platform connecting the small-scale farming community. It helps millions of African farmers meet, exchange solutions to their questions, and trade equipment and supplies. Farmers ask each other questions about agriculture and promptly receive content and ideas from other farmers worldwide through crowdsourcing. Natural language processing processes and understands the text-based messages farmers send in different regions. Machine learning matching algorithms consider farmers’ needs and expertise, identify the most suitable responses, and present them to the farmer who posed the question.
@@
Mercy Corps AgriFin 227 harnesses the power of AI and a global network of partners to transform agriculture for smallholder farmers. By utilizing state-of-the-art imagery, modeling, and analysis, AI enables precise management of crops and adaptation strategies against climate change impacts. AgriFin’s integration of digital technology and data, supported by a global network, empowers these farmers in a digitally interconnected world.
@@
In light of the global food crisis and the quest for sustainable solutions, the importance of AI-enhanced collective intelligence platforms becomes increasingly critical. These platforms can significantly contribute to addressing agricultural challenges and transitioning towards global sustainable food systems 228.
Limitations
The previous examples across various application domains demonstrate that CI and AI leverage the power of crowdsourced data, collaborative problem-solving, and advanced data processing techniques to enhance decision-making, optimize processes, and improve outcomes. However, there are limitations and concerns associated with the AI-CI approach.
Scalability
AI-enhanced collective intelligence can scale up problem-solving efficiency for real-world challenges, but scalability remains a limitation. As the system grows, scaling the multilayer framework and ensuring its applicability across various domains and scales can be challenging. Integrating diverse fields provides a comprehensive approach but poses difficulties in coordination, consistency, and computational resources.
Bias
While AI can help mitigate certain human biases, they also have the possibility to introduce new ones. These biases can stem from various stages of AI development and deployment. AI systems learn from the data they are trained on, and biased training data can cause the AI to replicate these biases. Additionally, algorithmic design choices may inadvertently favor certain groups over others. Human-in-the-loop bias is another concern, where human inputs during training and feedback, such as biased labeling by annotators can be learned by the AI. Moreover, user interactions with AI systems can also influence and reinforce these biases. AI biases pose risks to fair and equitable outcomes in society, especially in sensitive areas such as healthcare and education. Bias in AI algorithms can lead to unfair treatment or misdiagnosis in healthcare, and biased educational tools can perpetuate existing inequalities. Understanding how AI biases and human biases interact, and how to avoid doubly biased decisions made by human-machine intelligence 229, are crucial areas for further study.
Explainability
Many AI models, particularly deep learning algorithms, operate as “black boxes” 230, making understanding how they arrive at specific decisions or recommendations difficult. This opacity can hinder trust and acceptance among users, especially in high-stakes domains like healthcare and law. Explainability is essential for ensuring accountability, enabling users to understand and challenge AI decisions, and identifying and correcting biases within the system. Developing AI models that are both effective and interpretable requires ongoing research and innovation.
Challenges and outlooks
@@
The application areas of AI-enhanced collective intelligence continually expand as technology advances and new opportunities emerge. As introduced in the previous section, successful examples of AI-enhanced collective intelligence exist in various domains. However, challenges coexist with opportunities. When realizing human-AI collective intelligence, many aspects must be considered, such as human-machine communication, trust, crowd retention, technology design, and ethical issues 231. Here, we briefly discuss some of these challenges and call for further work on each aspect.
Communication
@@
Clear and effective communication between humans and AI is the engine for collective intelligence to emerge 168. Communication refers to the process of exchanging information between teammates 232. It is essential for team performance as it contributes to the development and maintenance of shared mental models and the successful execution of necessary team processes 232, 233.